Solve the following system of linear equations by graphing. (If the system has no solution, enter NONE for each answer. If the system is dependent, enter INFINITELY MANY for each answer.)
6x-2y=6
x-y=3
x = ?
y = ?
Please show me the steps to get the answer. Thanks

Answer by mananth(16946)

First generate points to draw each line.
6x-2y=6
when x= 0 y=-3 (0,-3)
when y =0, x=1 (1,0)
x-y=3
x=0 , y=-3 (0,-3)
y=0, x=3 9 (3,0)
You have the y & x intercept points of each line.

The lines intersect at (0,-3)